# Flux-balance laws for spinning bodies under the gravitational self-force

### v1

@@ II.2 Linear-in-spin motion
Equations (26) do not provide a complete set of equations for the 13 unknowns $\dot{\gamma}^{a}$, $p_{a}$, and $S^{ab}$: these are only 10 equations, and so we must supplement them with an additional choice in order to fix the location of the worldline. Such a choice is called a spin-supplementary condition, and we will adopt the Tulczyjew-Dixon condition [43, 4] for the remainder of this paper:
@@
Now, one thing to note about the motion described by Eqs. (28) is that it is not, in fact, equivalent to Eqs. (26) with the Tulczyjew-Dixon condition satisfied: instead, one only prove from Eq. (28) that
@@
As it simplifies the analysis, instead of considering the “true” motion in this paper, we will be considering Eq. (28) as given, and then apply $d^{a}=0$ as an initial condition, which by Eq. (30) implies that the Tulczyjew-Dixon condition holds for all time.
@@
where we consider the mass $m$ explicitly as a function on phase space. Hamilton’s equations also show that
$\dot{\gamma}^{a}=\frac{1}{m}g^{ab}p_{b}.$ (35)
For the spinning case, the situation is more subtle; in particular, if one were to use
@@
as coordinates on the 14-dimensional phase space, then the Hamiltonian remains the same as in Eq. (34), but the non-zero components of the Poisson bivector are given by [44]
@@
$\{S^{\alpha\beta},S^{\gamma\delta}\}$ $=2(g^{\alpha[\delta}S^{\gamma]\beta}-g^{\beta[\delta}S^{\gamma]\alpha}).$ (37d)
This means that $\Pi^{AB}$ is metric-dependent, and as we will see below, this implies that $\delta\Pi^{AB}\neq 0$, which significantly complicates the analysis.
Instead, following [45], we define better coordinates for this problem by first introducing the orthonormal tetrad $\boldsymbol{e}_{\hat{\alpha}}$, defined by
@@
As can be readily shown, the background motion (to first order in spin) can be described by Hamilton’s equations (see, for example, [46, 47, 39]), where $H$ is defined as in Eq. (34), but with $m$ defined as a function of $\pi_{\alpha}$ and $\hat{S}^{\hat{\alpha}\hat{\beta}}$ through a variable which we call $p_{a}(X)$:
@@
$p_{a}(X)\equiv\pi_{a}+\frac{1}{2}(\omega_{\hat{\alpha}\hat{\beta}})_{a}\hat{S}^{\hat{\alpha}\hat{\beta}},$ (45)
and the only non-trivial components of the Poisson bivector are determined by
@@
One final thing to note about our description of the background motion through the Hamiltonian (44) is that Hamilton’s equations only imply that Eq. (28) holds, not Eq. (26) with the Tulczyjew-Dixon condition. That is, this is a Hamiltonian for the “fictitious” problem that does not correspond to the physical motion, and the physical motion is only recovered when the initial data is such that Eq. (27) holds. This will be sufficient for our purposes, and (as we will show below in Sec. III.2) is necessary for our results to hold in the form they do.
@@
where $\breve{p}_{a}$ is defined analogously to how it is defined in Eq. (45), except by using $\breve{g}_{ab}(\bar{X};\varepsilon)$ instead of $g_{ab}$. As mentioned above, since the components of $\Pi^{AB}$ in these coordinates are independent of $g_{ab}$, there is no modification of $\Pi^{AB}$ in this gauge.
@@ III Field equations
Now, suppose that $\Phi_{\mathscr{A}}$ is a function of a single parameter $\lambda$. It then follows that one can define, from $\boldsymbol{Q}$, a local, linear functional $\boldsymbol{Q}^{\prime}$ by writing
@@
where we have dropped the functional dependence on $\Phi_{\mathscr{A}}$ on both sides, but we have explicitly included the dependence of $\boldsymbol{Q}^{\prime}$ on $\lambda$. When we set $\lambda=0$, we write the partial derivative as a variation $\delta$ [as in Eq. (13)] and omit the dependence on $\lambda$ from $\boldsymbol{Q}^{\prime}$:
@@
Note that $\boldsymbol{Q}^{\prime}(\lambda)$ is still a (potentially) nonlinear functional of $\Phi_{\mathscr{A}}$, although we are neglecting this dependence in these equations. To expand this dependence, and therefore get more information about the nonlinear dependence of $\boldsymbol{Q}$ on $\boldsymbol{\Phi}$, we apply the following procedure. First, consider a local, $n$-linear functional, which we generally write as $\boldsymbol{\mathcal{Q}}(\lambda)$, applied with the arguments $\Phi^{1}_{\mathscr{A}}(\lambda)$, …, $\Phi^{n}_{\mathscr{A}}(\lambda)$. Much like $\boldsymbol{Q}$, we can take a derivative of this functional, applied to these arguments, in order to define $\boldsymbol{\mathcal{Q}}^{\prime}(\lambda)$:
$\begin{split}\boldsymbol{\mathcal{Q}}^{\prime}(\lambda)&\left\{\frac{\partial\boldsymbol{\Phi}}{\partial\lambda},\boldsymbol{\Phi}^{1}(\lambda),\ldots,\boldsymbol{\Phi}^{n}(\lambda)\right\}\\ &\equiv\frac{\partial\boldsymbol{\mathcal{Q}}(\lambda)\{\boldsymbol{\Phi}^{1}(\lambda),\ldots,\boldsymbol{\Phi}^{n}(\lambda)\}}{\partial\lambda}\\ &\hskip 11.00008pt-\sum_{i=1}^{n}\boldsymbol{\mathcal{Q}}(\lambda)\left\{\boldsymbol{\Phi}^{1}(\lambda),\ldots,\frac{\partial\boldsymbol{\Phi}^{i}}{\partial\lambda},\ldots,\boldsymbol{\Phi}^{n}(\lambda)\right\}.\end{split}$ (53)
Unlike above, we explicitly eliminate any derivatives of the arguments using the sum on the right-hand side; in other words, we are taking the derivative with respect to $\lambda$, with all of the arguments of $\boldsymbol{\mathcal{Q}}$ fixed. In the case where we set $\lambda=0$, this becomes
$\begin{split}\boldsymbol{\mathcal{Q}}^{\prime}\{\delta\boldsymbol{\Phi},\boldsymbol{\Phi}^{1},\ldots,\boldsymbol{\Phi}^{n}\}&=\delta\boldsymbol{\mathcal{Q}}\{\boldsymbol{\Phi}^{1},\ldots,\boldsymbol{\Phi}^{n}\}\\ &\hskip 11.00008pt-\sum_{i=1}^{n}\boldsymbol{\mathcal{Q}}\{\boldsymbol{\Phi}^{1},\ldots,\delta\boldsymbol{\Phi}^{i},\ldots,\boldsymbol{\Phi}^{n}\}.\end{split}$ (54)
For a concrete example, we can use this procedure to define $\boldsymbol{Q}^{\prime\prime}$ by
@@
where the second line makes it apparent that $\boldsymbol{Q}^{\prime\prime}$, which is essentially the “quadratic part” of $\boldsymbol{Q}$, is symmetric in its two arguments.
With this notation established, in the remainder of this section we first describe a local, bilinear current $\boldsymbol{\omega}$, the symplectic current, which is conserved when the source-free, perturbative equations of motion are satisfied. We then turn to the discussion of the perturbative field equations from which the effective metric $\breve{g}_{ab}[\gamma;\varepsilon]$ is constructed in the first-order gravitational self-force, using the retarded, singular, and regular metric perturbations $h_{ab}$, $h^{S}_{ab}$, and $h^{R}_{ab}$, respectively. We conclude with a discussion of an important property relating the stress-energy tensor from which these metric perturbations are constructed and the Hamiltonian which appears in the self-force equations of motion.
@@ III.1 Symplectic currents
The definition of the symplectic current begins with considering the Lagrangian formulation of the theory in question; while in this paper we only consider general relativity, the discussion here can be generalized to an arbitrary field theory $\Phi_{\mathscr{A}}$. Unlike usual treatments of the Lagrangian formulation, we adopt the approach of [53, 54] in working entirely locally: instead of using an action, we consider only the variation of the integrand which appears in the action, which is naturally given by a Lagrangian four-form $\boldsymbol{L}$. This description has two advantages over using the usual action approach: first, one does not need to introduce a compensating boundary term in theories such as general relativity where the Lagrangian depends on second derivatives of the field. Second, the boundary term arising in the variation, which is typically neglected at various points in the action approach, appears far more prominently.
@@
Note, however, that as it stands, there are two (related) issues with using the symplectic product in order to define a flux at infinity. First, it is a bilinear, antisymmetric current, which requires that we have two different field perturbations on the background spacetime, each satisfying the equations of motion, in order to construct a conserved current. Second, physically relevant fluxes are typically quadratic in the field. However, if we were to have an operator, which we denote by $\mathcal{D}_{\mathscr{A}}{}^{\mathscr{B}}$, which maps the space of solutions to the vacuum equations of motion to itself, we could resolve both of these issues and use such a map to define a quadratic current from a single perturbation [54]. Such an operator is called a symmetry operator [55].
@@ III.2 Self-force
where, from Eq. (62) (and the fact that we are in a vacuum background spacetime), $(\boldsymbol{E}^{\prime})^{ab}$ is a differential-form version of the linearized Einstein operator, and $\boldsymbol{T}^{ab}[\gamma]$ is a distributional source that is given below in Eq. (74). In particular, it is a vacuum solution off of $\gamma$. In a convex normal neighborhood of $\gamma$, one can split the retarded field $h_{ab}$ into two fields, the singular field $h^{S}_{ab}$, which is the part which is not a vacuum solution, and in fact blows up, on $\gamma$:
@@
While it is unimportant for the discussion in this paper, one can operationally compute these fields as follows: first, up to a given order in distance from the source, $h^{S}_{ab}$ has a known analytic form (given, for example, in [51]). After solving for the full retarded solution $h_{ab}$, one can then simply subtract off $h^{S}_{ab}$ in order to obtain $h^{R}_{ab}$. Typically, this is only done at some truncated order in distance to the particle, but at sufficiently high order that the value of $h^{R}_{ab}$ and its first derivative can be recovered exactly. At second order, this discussion becomes more complicated, see for example [64].
@@
As such, this stress-energy tensor is supported entirely on $\gamma$. Note that, upon taking a derivative $\nabla_{b^{\prime}}$ of Eq. (71), we find
@@
Using Eq. (77), and comparing with Eq. (74), we find that the Hamiltonian alignment condition holds, with $C_{g}=-1/2$. To linear order in spin, we have that
@@
where in the first equality we have used the fact that ˘$\breve{\nabla}_{a}$ only depends on $\nabla_{a}h^{R}_{bc}$ to linear order. As such, the second term in  (79) doesn’t contribute, and so
@@
$(\partial H)^{abc}(X)=\frac{g^{cd}p_{c}(\partial\omega_{\hat{\alpha}\hat{\beta}})_{d}{}^{abc}\hat{S}^{\hat{\alpha}\hat{\beta}}}{2\mu(X)}.$ (83)
@@
By comparing Eqs. (82) and (87) to Eq. (74) and (77), we see that, even in the linear-in-spin case, the Hamiltonian alignment condition holds, with $C_{g}=-1/2$.
@@ IV General flux-balance laws
As in [33], we can consider the operation given by taking a derivative with respect to $X$, which is a symmetry operator which we denote by $\nabla_{A}$.
Using this symmetry operator, we can now define a flux integral over some hypersurface, which we generically denote by $\Sigma(\tau,\Delta\tau)$. We assume that this hypersurface is a tube surrounding the worldline $\gamma$, with boundaries which we denote by $\partial\Sigma_{\pm}$. These surfaces $\partial\Sigma_{\pm}$ are also the boundaries of hypersurfaces $\Sigma_{\pm}$ which intersect $\gamma$ at $\gamma(\tau_{\pm})$, where $\tau\pm\equiv\tau\pm\Delta\tau/2$. The two hypersurfaces we will be integrating our fluxes over will consist of an outer boundary $\mathcal{B}(\tau,\Delta\tau)$, which we will later assume we can extend down to the event horizon $\mathscr{H}$ and up to null infinity $\mathscr{I}$, and a worldtube $\mathcal{W}(\tau,\Delta\tau;r)$ of some proper radius $r$. Note that $\mathcal{W}(\tau,\Delta\tau;r)$ needs to be within the convex normal neighborhood of $\gamma$. As a visual aide, see Fig. 1.
@@
where $X$, the point at which $\nabla_{A}$ is applied, is given by $\Gamma(\tau)$, the integrand is evaluated at some point $x^{\prime}$ on the surface $\Sigma(\tau,\Delta\tau)$, and where $C_{g}$ is the constant (in gravity equal to $-1/2$) appearing in the Hamiltonian alignment condition in Eq. (76).
@@
$\mathcal{F}^{\mathcal{B}}_{A}\{\boldsymbol{h},\boldsymbol{h}\}=\mathcal{F}^{\mathcal{W}(r)}_{A}\{\boldsymbol{h},\boldsymbol{h}\},$ (90)
@@
$\left\langle\delta\dot{\Gamma}^{A}(\tau^{\prime})\right\rangle=\Pi^{AB}\bigg{[}\frac{1}{2}\mathcal{F}^{\mathcal{B}}_{B}\{\boldsymbol{h},\boldsymbol{h}\}+\nabla_{B}\langle\mathcal{H}(X,\tau^{\prime})\rangle_{\tau^{\prime}}\bigg{]},$ (91)
where $\mathcal{H}$, defined by
$\mathcal{H}(X,\tau^{\prime})\equiv\frac{1}{2}H[\Upsilon_{\tau^{\prime}-\tau}(X),X],$ (92)
is what we call the “coincidence Hamiltonian”. This object is similar to the perturbed Hamiltonian describing the conservative dynamics which appears in, for example, [38, 39, 40], although their Hamiltonian is constructed only from the time-symmetric part of the regular field $h^{R}_{ab}$, and the two points are assumed to be the same, with $\tau^{\prime}=\tau$.

### v4

@@ II.2 Linear-in-spin motion
Equations (26) do not provide a complete set of equations for the 13 unknowns $\dot{\gamma}^{a}$, $p_{a}$, and $S^{ab}$: these are only 10 equations, and so we must supplement them with an additional choice in order to fix the location of the worldline. Such a choice is called a spin-supplementary condition, and we will adopt the Tulczyjew-Dixon condition [55, 4] for the remainder of this paper: [^8] As we are explicitly excluding the null case, we are free to ignore the known pathologies of this spin-supplementary condition in this case (see, for example, the discussion in [56, 57] and pg. 70 of [58]).
@@
Now, one thing to note about the motion described by Eqs. (28) is that it is not, in fact, equivalent to Eqs. (26) with the Tulczyjew-Dixon condition satisfied: instead, one can only prove from Eq. (28) that
@@
Instead of considering the “true”, restricted phase space for spinning bodies in this paper, we will be considering Eq. (28) as given, and then apply $d^{a}=0$ as an initial condition, which by Eq. (30) implies that the Tulczyjew-Dixon condition holds for all time. The Hamiltonian which we consider below is therefore for a “fictitious” problem that does not correspond to the physical motion, and the physical motion is only recovered when the initial data are such that Eq. (27) holds. This will be sufficient for our purposes, and (as we will show below in Sec. III.2) it is necessary for our results to hold in the form they do.
@@
where we consider the mass $m$, as defined by Eq. (29), explicitly as a function on phase space. Hamilton’s equations also show that
$\dot{\gamma}^{a}=\frac{1}{m}g^{ab}p_{b},$ (35)
directly enforcing the normalization of $\dot{\gamma}^{a}$.
For the spinning case, the situation is more subtle. First, suppose that one uses
@@
as coordinates on the 14-dimensional phase space. If one uses the Poisson bivector given by
@@
$\{S^{\alpha\beta},S^{\gamma\delta}\}$ $=2(g^{\alpha[\delta}S^{\gamma]\beta}-g^{\beta[\delta}S^{\gamma]\alpha})$ (37d)
(a natural choice appearing throughout the literature [59, 60, 61]), then the Hamiltonian for the MPD equations with the Tulczjew-Dixon spin supplementary condition is still given by Eq. (34) (see [62] for an extensive discussion of the relationship between the choice of spin-supplementary condition and Hamiltonian). This means that $\Pi^{AB}$ is metric-dependent, and as we will see below, this implies that $\delta\Pi^{AB}\neq 0$, which significantly complicates the analysis.
Instead, following [60, 63], we define better coordinates for this problem by first introducing the orthonormal tetrad $\boldsymbol{e}_{\hat{\alpha}}$, defined by
@@
Since, as mentioned above, $H$ in the original coordinates is given by Eq. (34), it follows that in these good coordinates it is given by the same equation, but with $m$ defined as a function of $\pi_{\alpha}$ and $\hat{S}^{\hat{\alpha}\hat{\beta}}$ through a variable which we call $p_{a}(X)$:
@@
$p_{a}(X)\equiv\pi_{a}+\frac{1}{2}(\omega_{\hat{\alpha}\hat{\beta}})_{a}\hat{S}^{\hat{\alpha}\hat{\beta}}.$ (45)
What has changed is that the non-zero components of the Poisson bivector are now determined by [60, 63]
@@
where $\breve{p}_{a}$ is defined analogously to how it is defined in Eq. (45), except by using a $(\breve{e}_{\hat{\alpha}})^{a}(\bar{X};\varepsilon)$ and $\breve{\omega}_{\hat{\alpha}\hat{\beta}}(\bar{X};\varepsilon)$ that are defined from $\breve{g}_{ab}(\bar{X};\varepsilon)$ and its compatible covariant derivative, instead of $g_{ab}$, using Eqs. (38) and (41), respectively. Note that $(e^{\hat{\alpha}})_{a}$ does not appear in Eq. (45), and so $(\breve{e}^{\hat{\alpha}})_{a}(\bar{X};\varepsilon)$ does not arise in this equation. As mentioned above, since the components of $\Pi^{AB}$ in these coordinates are independent of $g_{ab}$, there is no modification of $\Pi^{AB}$ in this gauge to an analogous $\breve{\Pi}^{AB}$.
@@ III Field equations
Now, suppose that $\Phi_{\mathscr{A}}$ is a function of a single parameter $\lambda$. It then follows that one can define, from $\boldsymbol{Q}$, a local, linear functional $\boldsymbol{Q}^{\prime}(\lambda)$ by writing
@@
where we have dropped the functional dependence on $\Phi_{\mathscr{A}}$ on both sides, but we have explicitly included the dependence on $\lambda$. When we set $\lambda=0$, we write the partial derivative as a variation $\delta$ [as in Eq. (13)] and omit the dependence on $\lambda$ from $\boldsymbol{Q}^{\prime}$:
@@
While $\boldsymbol{Q}^{\prime}(\lambda)$ reflects the linear dependence of $\boldsymbol{Q}$ on $\boldsymbol{\Phi}$, the quadratic dependence of $\boldsymbol{Q}$ on $\boldsymbol{\Phi}$ will also be occasionally be relevant. This dependence can be recovered from the remaining dependence of $\boldsymbol{Q}^{\prime}(\lambda)$ on $\Phi_{\mathscr{A}}$, which we have neglected in the equations above. This can be seen as follows: for any local functional $\boldsymbol{\mathcal{Q}}(\lambda)$, which depends nonlinearly on $\Phi_{\mathscr{A}}(\lambda)$ and linearly on some $\phi_{\mathscr{A}}(\lambda)$, we take a derivative with respect to $\lambda$ to define $\boldsymbol{\mathcal{Q}}^{\prime}(\lambda)$:
$\boldsymbol{\mathcal{Q}}^{\prime}(\lambda)\left\{\frac{\partial\boldsymbol{\Phi}}{\partial\lambda},\boldsymbol{\phi}(\lambda)\right\}\equiv\frac{\partial\boldsymbol{\mathcal{Q}}(\lambda)\{\boldsymbol{\phi}(\lambda)\}}{\partial\lambda}-\boldsymbol{\mathcal{Q}}(\lambda)\left\{\frac{\partial\boldsymbol{\phi}}{\partial\lambda}\right\}.$ (53)
Unlike above, we explicitly eliminate any derivatives of the arguments using the second term on the right-hand side; in other words, we are taking the derivative with respect to $\lambda$, with the argument of $\boldsymbol{\mathcal{Q}}$ fixed. In the case where we set $\lambda=0$, this becomes
$\boldsymbol{\mathcal{Q}}^{\prime}\{\delta\boldsymbol{\Phi},\boldsymbol{\phi}\}=\delta\boldsymbol{\mathcal{Q}}\{\boldsymbol{\phi}\}-\boldsymbol{\mathcal{Q}}\{\delta\boldsymbol{\phi}\}.$ (54)
We can now use this procedure to define $\boldsymbol{Q}^{\prime\prime}$ by
@@
where the second line makes it apparent that $\boldsymbol{Q}^{\prime\prime}$, which is essentially the “quadratic part” of $\boldsymbol{Q}$, is symmetric in its two arguments. It follows from these definitions that
$\begin{split}\boldsymbol{Q}(\lambda)&=\boldsymbol{Q}+\lambda\boldsymbol{Q}^{\prime}\{\delta\boldsymbol{\Phi}\}+\frac{\lambda^{2}}{2}\left(\boldsymbol{Q}^{\prime}\{\delta^{2}\boldsymbol{\Phi}\}+\boldsymbol{Q}^{\prime\prime}\{\delta\boldsymbol{\Phi},\delta\boldsymbol{\Phi}\}\right)\\ &\hskip 11.00008pt+O(\lambda^{3}).\end{split}$ (56)
With this notation established, in the remainder of this section we first describe a local, bilinear current $\boldsymbol{\omega}$, the symplectic current, which is conserved when the source-free, perturbative equations of motion are satisfied. We then turn to the discussion of the perturbative field equations from which the effective metric $\breve{g}_{ab}[\gamma;\varepsilon]$ is constructed in the first-order gravitational self-force, using the retarded, singular, and regular metric perturbations $h_{ab}$, $h^{S}_{ab}$, and $h^{R}_{ab}$, respectively. We conclude with a discussion of an important property relating the stress-energy tensor from which these metric perturbations are constructed and the pseudo-Hamiltonian which appears in the self-force equations of motion.
@@ III.1 Symplectic currents
The definition of the symplectic current begins with considering the Lagrangian formulation of the theory in question; while in this paper we only consider general relativity, the discussion here can be generalized to a theory for an arbitrary field $\Phi_{\mathscr{A}}$. Unlike usual treatments of the Lagrangian formulation, we adopt the approach of [69, 70] in working entirely locally: instead of using an action, we consider only the variation of the integrand which appears in the action, which is naturally given by a Lagrangian four-form $\boldsymbol{L}$. This description has two advantages over using the usual action approach: first, one does not need to introduce a compensating boundary term in theories such as general relativity where the Lagrangian depends on second derivatives of the field. Second, the boundary term arising in the variation, which is typically neglected at various points in the action approach, appears far more prominently.
@@
Note, however, that there is an issue with using the symplectic product in order to define a flux at infinity: it is a bilinear, antisymmetric current, which requires that we have two different field perturbations on the background spacetime, each satisfying the equations of motion, in order to construct a conserved current. On the other hand, physically relevant fluxes are typically quadratic in the field. However, if we were to have an operator, which we denote by $\mathcal{D}_{\mathscr{A}}{}^{\mathscr{B}}$, which maps the space of solutions to the vacuum equations of motion to itself, we could resolve this issue by using such a map to define a quadratic current from a single perturbation [70]. Such an operator is called a symmetry operator [71].
@@ III.2 Self-force
where, from Eq. (63) (and the fact that we are in a vacuum background spacetime), $(\boldsymbol{E}^{\prime})^{ab}$ is a differential-form version of the linearized Einstein operator and $\boldsymbol{T}^{ab}[\gamma]$ is a distributional source that is given below in Eq. (76). In particular, $h_{ab}$ is a vacuum solution off of $\gamma$. In a convex normal neighborhood of $\gamma$, one can split the retarded field $h_{ab}$ into two fields, the singular field $h^{S}_{ab}$, which is the part which is not a vacuum solution, and in fact blows up, on $\gamma$:
@@
While it is unimportant for the discussion in this paper, one can operationally compute these fields as follows: first, up to a given order in distance from the source, $h^{S}_{ab}$ has a known analytic form (given, for example, in [67]). After solving for the full retarded solution $h_{ab}$, one can then simply subtract off $h^{S}_{ab}$ in order to obtain $h^{R}_{ab}$. Typically, this is only done at some truncated order in distance to the particle, but at sufficiently high order that the value of $h^{R}_{ab}$ and its first derivative can be recovered exactly. At second order, this discussion becomes more complicated; see for example [80].
@@
for any collections of indices $\mathscr{A}$ and $\mathscr{B}$. As such, this stress-energy tensor is supported entirely on $\gamma$. Moreover, the distribution $\boldsymbol{\delta}^{\mathscr{A}}{}_{\mathscr{A}^{\prime}}$ satisfies the property that
$\boldsymbol{\delta}^{\mathscr{A}\mathscr{B}}{}_{\mathscr{A}^{\prime}\mathscr{B}^{\prime}}=\boldsymbol{\delta}^{\mathscr{B}\mathscr{A}}{}_{\mathscr{B}^{\prime}\mathscr{A}^{\prime}}.$ (73)
Further, note that, upon taking a derivative $\nabla_{b^{\prime}}$ of Eq. (72), we find
@@
Using Eq. (79), and comparing with Eq. (76), we find that the Hamiltonian alignment condition holds, with $C_{g}=-1/2$. There are no terms in Eq. (79) beyond $n=0$, since the pseudo-Hamiltonian in this case does not depend upon derivatives of the metric perturbation. To linear order in spin, we have that
@@
where in the first equality we have used the fact that ˘$\breve{\nabla}_{a}$ only depends on $\nabla_{a}h^{R}_{bc}$ to linear order. As such, the second term in Eq. (81) doesn’t contribute, and so
@@
$(\partial H)^{abc}(X)=\frac{g^{cd}p_{c}(\partial\omega_{\hat{\alpha}\hat{\beta}})_{d}{}^{abc}\hat{S}^{\hat{\alpha}\hat{\beta}}}{2m(X)}.$ (85)
@@
By comparing Eqs. (84) and (89) to Eq. (76) and (79), we see that, even in the linear-in-spin case, the Hamiltonian alignment condition holds, with $C_{g}=-1/2$. We do not need to go beyond $n=1$ in Eq. (77) because there are no terms in the pseudo-Hamiltonian which depend on more than one derivative of the metric perturbation.
@@ IV General flux-balance laws
As in [41], we then consider the operation given by taking a derivative with respect to $X$, which (as before) we denote by $\nabla_{A}$. Note that derivatives of $h^{\cdots}_{a^{\prime}b^{\prime}}(X)$ with respect to $X$ and $x^{\prime}$ act independently since, considered as a bitensor at $X$ and $x^{\prime}$, $h^{\cdots}_{a^{\prime}b^{\prime}}(X)$ does not impose any relationship through its definition (unlike, say, $\Upsilon^{A^{\prime}}{}_{A}$). As such, $\nabla_{A}$ and $\nabla_{a^{\prime}}$ commute, and moreover $\nabla_{A}g_{a^{\prime}b^{\prime}}=0$ (the background metric is not a function of $X$), so $\nabla_{A}$ is a symmetry operator.
Using this symmetry operator, we can define a flux integral over some hypersurface, which we generically denote by $\Sigma(\tau,\Delta\tau)$. We assume that this hypersurface is a tube surrounding the worldline $\gamma$, with boundaries which we denote by $\partial\Sigma_{\pm}$. These surfaces $\partial\Sigma_{\pm}$ are also the boundaries of hypersurfaces $\Sigma_{\pm}$ which intersect $\gamma$ at $\gamma(\tau_{\pm})$, where $\tau\pm\equiv\tau\pm\Delta\tau/2$. The two hypersurfaces we will be integrating our fluxes over will consist of an outer boundary $\mathcal{B}(\tau,\Delta\tau)$, which we will later assume can be extended down to the event horizon $\mathscr{H}$ and out to null infinity $\mathscr{I}$, and a worldtube $\mathcal{W}(\tau,\Delta\tau;r)$ of some proper radius $r$. Note that $\mathcal{W}(\tau,\Delta\tau;r)$ needs to be within the convex normal neighborhood of $\gamma$. As a visual aid, see Fig. 1.
@@
where $X$, the point at which $\nabla_{A}$ is applied, is given by $\Gamma(\tau)$; the integrand is evaluated at some point $x^{\prime}$ on the surface $\Sigma(\tau,\Delta\tau)$; and $C_{g}$ is the constant (in gravity equal to $-1/2$) appearing in the Hamiltonian alignment condition in Eq. (78).
@@
$\mathcal{F}^{\mathcal{B}}_{A}\{\boldsymbol{h},\boldsymbol{h}\}=\mathcal{F}^{\mathcal{W}(r)}_{A}\{\boldsymbol{h},\boldsymbol{h}\}$ (92)
@@
$\begin{split}\left\langle\delta\dot{\Gamma}^{A}(\tau^{\prime})\right\rangle_{\tau^{\prime}}=\Pi^{AB}\bigg{\{}\frac{1}{2}\mathcal{F}^{\mathcal{B}}_{B}\{\boldsymbol{h},\boldsymbol{h}\}+\nabla_{B}\langle\mathcal{H}(X^{\prime})\rangle_{\tau^{\prime}}\bigg{\}},\end{split}$ (93)
where for brevity we write $X^{\prime}\equiv\Upsilon_{\tau^{\prime}-\tau}(X)$, and $\mathcal{H}$ is the Hamiltonian for the conservative dynamics described in detail in [49, 50, 51], and discussed in more detail in Sec. IV.2 below.